# Enhancing Peer Review in Astronomy: A Machine Learning and Optimization Approach to Reviewer Assignments for ALMA

### v1

@@ 4.2 Reviewer expertise
The PI of a proposal must specify one person from the proposal team that will serve as a reviewer in the review process. Most PIs serve as the reviewer on behalf of their proposals, but in about 12% of cases, a co-Investigator (coI) is selected (Donovan Meyer et al., 2022). The reviewer must be a member of the proposal team and registered within the ALMA system.
In previous applications of machine learning to proposal assignments at astronomical observatories (Kerzendorf, 2019; Kerzendorf et al., 2020; Strolger et al., 2023), the reviewer expertise has been inferred based on published abstracts or papers that are available in ADS. This approach has the advantage in that it contains an extensive record of a reviewer’s publication history and therefore expertise. While the publication record of most authors can be inferred accurately from their last name and first initial (Milojević, 2013), erroneous matches will inevitably occur. With over 1000 reviewers participating in the review process, it is not practical to perform a manual inspection of the search results to identifying any problematic identifications. While an Open Researcher and Contributor ID (ORCID) can be used to identify authors uniquely, not all papers are tagged with an ORCID, [^4] https://orcid.org and ALMA does not currently collect ORCIDs from its users. Also, ALMA allows students who are the PI of proposals to serve as reviewers by specifying a mentor, and students may not yet have a publication record.
Instead of using the publication record, ALMA used the proposal history of the reviewers, including the current proposal cycle, to establish their scientific expertise. The main advantage of this approach is that reviewers can be uniquely identified by their ALMA user ID, and by design of the review process, each reviewer is guaranteed to be an investigator for at least one proposal. In addition, each proposal is tagged by the proposal PI with a scientific category and keywords that provides additional information on the topic of the proposal. As discussed below, these keywords can help identify which proposals should be used to determine the reviewer’s expertise and to prevent proposal assignments in disparate scientific categories (see Section 4.3). A downside to using the proposal text is that the proposal history will not capture the full expertise of a reviewer, especially for senior astronomers who are new users to ALMA and may have a substantial publication history.
@@
The similarity score $S(\vec{p_{i}},\vec{r_{k}})$ ranges from 0 to 1, where 0 indicates no alignment between the reviewer’s expertise and the proposal topic, and 1 indicates a perfect match. While the absolute value of the similarity score may not be crucial, the relative values should provide a meaningful comparison of how well each proposal aligns with the reviewer’s expertise.
This general approach works well for most reviewers, but the implementation also accounted for those with expertise or proposal submissions across multiple categories. Special considerations were also made for categories and keywords with limited available reviewers. The Appendix provides further details on how reviewer expertise was calculated to address these specific situations.
@@ 4.3 Proposal assignments
Before assigning proposals, similarities are set to a negative number if a proposal assignment is not permitted. There are two main reasons to restrict proposal assignments in this manner. First, potential assignments are flagged if the reviewer has a known conflict of interest on the proposal. Conflicts of interest are defined if the reviewer is a PI or coI on the potential proposal assignment proposal, the PI is deemed a close collaborator based on the proposal history (see Donovan Meyer et al., 2022), or the reviewer identified a PI or coI as a close-collaborator. In these cases, the similarities are set to a value of $-100$, which effectively prohibits the assignment of the conflicted proposal. The second case is if the potential assigned proposal is from a distinct scientific category compared to the reviewer’s expertise (see Section 4.2). While such proposals will typically have low similarities, this additional measure reduces the risk of a false positive match. In practice, the new similarity is set as $-1.01+S(\vec{p_{i}},\vec{r_{k}})$. While a negative number does not preclude a proposal assignment, it makes it less likely and can be checked after the assignments are made.
@@
Once the similarity matrix is established, proposals are assigned to reviewers. The goal is to define a metric that measures the success of these assignments and optimize it during the assignment process. One common approach is to maximize the total similarity scores between proposals and reviewers, which increases overall expertise in the review process. However, this method may lead to unfairness if certain types of proposals are disproportionately assigned to less-qualified reviewers. ALMA instead adopted the approach used in PeerReview4All, which optimizes a “leximin fairness” metric (Stelmakh et al., 2021a, code available at https://github.com/niharshah/peerreview4all). This method optimizes the similarity for the most “disadvantaged” proposals; i.e., proposals that have the least overall expertise among the reviewers. Once the most disadvantaged proposal’s assignment is optimized, the algorithm moves to the next most disadvantaged, and so on. Proposals in fields with many qualified reviewers are assigned last, as they have more reviewer options available. The metric used by PeerReview4All is solved using the gurobi python optimizer (Gurobi Optimization, LLC, 2023).
@@
The PeerReview4All code was adapted to handle special cases. Student reviewers specify a mentor, who may also serve as a reviewer for other proposals. To ensure fairness, we required that the proposal sets for students and mentors be disjoint. After each round of assignments, the similarity matrix was updated to mark as conflicted any proposals previously assigned to a reviewer or their mentor. Additionally, conflicts were updated for reviewers assigned multiple sets, ensuring they were not assigned the same proposal twice.
@@
To prevent “niche” proposals from repeatedly being reviewed by the same group, further restrictions were imposed on the number of proposals a reviewer could be assigned within the same topic. These restrictions specifically applied to proposals related to Very Long Baseline Interferometry, Target of Opportunity, and solar studies.
5 Evaluation
@@
Given the framework for computing the similarity between proposals and reviewer’s expertise described in Section 4, we now evaluate if the new process improved the proposal assignments to the reviewers compared to previous cycles. Recall that the ALMA reviewing cycle where the new assignment approach was implemented is Cycle 10, and the two previous cycles we use for backtesting are Cycle 8 and Cycle 9.
5.1 Backtesting similarity computation on Cycle 8 and Cycle 9 data
@@
We backtest the effectiveness of the similarity scores (calculated as described in Section 4.2) by using data from previous review cycles. Specifically, we examine whether the computed similarities between a reviewer’s expertise and the proposal topics accurately reflect the reviewer’s expertise. During the distributed peer review process used by ALMA in Cycles 8 and 9, reviewers were asked to self-assess their expertise on each assigned proposal by indicating if the assigned proposal is in their field of expertise (i.e., the review considers themselves an “expert”), they are somewhat knowledgeable, or they have little to no knowledge of the topic. If our similarity scores are a reliable measure of expertise, we would expect that reviewers who self-identify as experts in a proposal topic will, on average, have higher similarity scores compared to those who report having little or no knowledge of the topic.
@@
Figure 3: Cumulative distributions of the similarities between the reviewer’s expertise and the topic of the assigned proposal in Cycle 8 (left) and Cycle 9 (right), when categories and keywords were used to assigned proposals. The similarities were computed retroactively to backtest the machine learning algorithm. The results are shown for proposal assignments where reviewers indicated they are experts (solid curve), had some knowledge of the proposal (dashed), and had little/no knowledge of the proposal (dotted). Similarities tend to be higher when reviewers indicate expertise in the proposal, suggesting that these similarities have predictive value for assessing the suitability of reviewer assignments.
5.2 Comparing Cycle 10 and Cycle 9 assignments, given computed similarities
In this section, we analyze the role of the optimization procedure for computing the assignments (Section 4.3) given the computed similarities. In Cycle 9, proposals were assigned to reviewers based on the proposal category and keywords (Donovan Meyer et al., 2022); i.e., the “old” algorithm. In Cycle 10, we used the similarities and assignment algorithm described in Section 4; i.e., the “new” algorithm. In order to make a fair comparison, we recompute similarities for Cycle 9 using the software developed for Cycle 10. In the analysis in this subsection, we use these similarities. In both cycles, a small percentage of the initial assignments were reassigned after reviewers declared conflicts of interest. [^5] The percentage of assignments where reviewers declared a conflict of interest increased in Cycle 10 to 2.4% from 1.4% in Cycle 9, as more reviewers declared the assignments conflicted scientifically with their own proposal when the machine learning algorithm was adopted.
Figure 4 shows a histogram of the similarities in the actual proposals assignments for Cycle 9 (blue histogram), and Cycle 10 (orange histogram) which used the optimization method of Section 4.3. We find a substantial increase in the similarity between the reviewer expertise and the assigned proposals compared to the previous algorithm, with the median similarity increase to 0.71 in Cycle 10 from 0.20 in Cycle 9.
@@ 5.3 Cycle 10 reviewer survey
The increased similarity of the proposal assignments was also reflected in the reviewer survey, where reviewers provided a self-assessment of their expertise on each proposal assignment. Figure 5 shows a histogram of the results from Cycles 8, 9 and 10. With the new assignment process adopted in Cycle 10, the percentage of reviewers indicating they were an expert in the assigned proposal increase to 65% in Cycle 10 from 45% in Cycle 9. Similarly, the percentage of assignments where the reviewers indicated they had little or no knowledge on the proposal declined to 5% in Cycle 10 from 10% in Cycle 9.
@@ 5.4 Topological accuracy of the proposal rankings
In this section, we evaluate a key tradeoff in expertise-based reviewer assignments: balancing the expertise of assigned reviewers with the topology of the assignments. Reviewer expertise is crucial for evaluating research proposals, prompting strategies that prioritize maximizing relevant expertise. However, strictly focusing on expertise can lead to the isolation of certain proposal groups, especially those within the same subtopic. For example, imagine a scenario where half of the reviewers are exclusively assigned to review one group of proposals, while the other half reviews the rest. This division hampers direct comparison between the two groups, limiting the ability to rank proposals effectively across groups.
@@
To estimate the relative accuracy of the actual assignments in Cycle 10, we set the error in the actual assignments to unity. We then ran simulations where the reviewers were assigned proposals at random across any scientific category, subject to the constraint that each reviewer is assigned 10 proposals and each proposal is assigned to 10 reviewers. The random assignment strategy is known to be optimal (Shah et al., 2016) in terms of topology when all reviewers have identical expertise. We find that the relative uncertainty is 0.86 for the random assignment. Thus, optimizing assignments for expertise increases the uncertainty in final outcomes by only 14% based on topology of the proposal assignments.
6 Discussion and conclusions
The implementation of the new assignment process in ALMA Cycle 10, leveraging machine learning (Blei et al., 2003) and optimization (Stelmakh et al., 2021a) algorithms, has successfully met the high-level objectives set for the system. The similarity measures between the proposal topics and reviewer’s expertise computed from their proposal history are correlated with the reviewer’s self-reported expertise to demonstrate the similarity measures have predictive value. With the new optimization algorithm, the average similarity is significantly higher than in previous cycles with the old algorithm. The success of the algorithm is confirmed in the Cycle 10 reviewer survey, which showed that reviewers reported a higher level of expertise for the assignment than the previous two cycles. Finally, the initial assignments were run without any manual adjustments, resulting in significant time savings in handling the proposals and indicating the robustness of the process.
Our rigorous evaluation demonstrates the potential for widespread adoption of AI-driven reviewer assignments in both proposal and paper reviewing, offering substantial improvements. This methodology is especially useful for scientific disciplines that continue to struggle with manual assignment processes, where scalability and growing complexity demand more sophisticated, automated approaches.
6.1 Limitations
We now discuss some limitations of our study. Firstly, a portion of our findings is based on reviewers’ self-reported expertise regarding the proposals they assessed. Such self-reports are susceptible to various biases, suggesting that future research should incorporate additional evaluative metrics. Secondly, our analysis largely compares the current cycle to previous ones. There is a possibility that other significant changes, such as shifts in the types of proposals or the reviewer demographic, could have influenced the results. To counteract potential confounding factors, we conducted an analysis using the current algorithms on data from previous cycles, which affirmed the consistency of our results. Additionally, the firsthand experience of the first two authors, who have overseen this review process for several cycles, indicates no major changes during this period.

### v2

@@ 2.2 Reviewer expertise
In previous applications of machine learning to proposal assignments at astronomical observatories (Kerzendorf, 2019; Kerzendorf et al., 2020; Strolger et al., 2023), the reviewer’s expertise has been inferred based on published abstracts or papers available in ADS. This approach has the advantage that it contains an extensive record of a reviewer’s publication history. While the publication record of most authors can be inferred accurately from their last name and first initial (Milojević, 2013), erroneous matches will inevitably occur. With over 1000 reviewers participating in the review process, it is not practical to perform a manual inspection of the search results to identify any problematic matches. While an Open Researcher and Contributor ID (ORCID) can be used to identify authors uniquely, not all papers are tagged with an ORCID, [^3] https://orcid.org and ALMA does not currently collect ORCIDs from its users. Also, ALMA allows students who are the PI of proposals to serve as reviewers by specifying a mentor, and students may not yet have a publication record.
Instead of using the publication record, we used the proposal history of the reviewers, including the current proposal cycle, to establish their scientific expertise. The inherent assumption is that a reviewer will have expertise in the science topics for proposals in which they are the PI or a co-investigator (coI). The main advantage of this approach is that reviewers are uniquely identified by their ALMA user ID, and, by design, each reviewer is guaranteed to be an investigator on at least one proposal. In addition, each proposal is tagged by the proposal PI with a scientific category and keywords that provide additional information on the topic of the proposal. As discussed below, these keywords can help identify which proposals should be used to determine the reviewer’s expertise and to prevent proposal assignments in disparate scientific categories (see Section 2.3). A downside to using the proposal text is that the proposal history will not capture the full expertise of a reviewer, especially for senior astronomers who are new users to ALMA and may have a substantial publication history.
@@
The similarity score $S(\vec{p_{i}},\vec{r_{k}})$ ranges from 0 to 1, where 0 indicates no alignment between the reviewer’s expertise and the proposal topic, and 1 indicates a perfect match. While the absolute values of the similarity scores may be less critical, their relative values are meaningful for comparing how well each proposal aligns with a reviewer’s expertise.
This general approach works well for most reviewers, but the implementation also accounted for those with expertise or proposal submissions in multiple categories. Special considerations were also made for categories and keywords with limited reviewers available. Appendix A provides further details on how reviewer expertise was calculated to address these specific situations.
@@ 2.3 Proposal assignments
Before assigning proposals, similarities are set to a negative number if a proposal assignment is not permitted. There are two main reasons for restricting proposal assignments in this manner. First, potential assignments are flagged if the reviewer has a known conflict of interest on the proposal. Conflicts of interest are defined if the reviewer is a PI or coI on the potential proposal assignment proposal, the PI is deemed a close collaborator based on the proposal history (see Donovan Meyer et al., 2022), or the reviewer identified a PI or coI as a close collaborator. In these cases, the similarities are set to a value of $-100$, which effectively prohibits the assignment of the conflicted proposal. The second case is if the potential assigned proposal is from a distinct scientific category compared to the reviewer’s expertise (see Section 2.2). While such proposals will typically have low similarities, this additional measure reduces the risk of a false positive match. In practice, the new similarity is set as $-1.01+S(\vec{p_{i}},\vec{r_{k}})$. While a negative number does not preclude a proposal assignment, it makes it less likely and can be checked after the assignments are made.
@@
Once the similarity matrix is established, proposals are assigned to reviewers. The goal is to define a metric that measures the success of these assignments and optimize it during the assignment process. One common approach is to maximize the total similarity scores between proposals and reviewers, which increases overall expertise in the review process. However, this method may lead to unfairness if certain types of proposals are disproportionately assigned to less-qualified reviewers. ALMA instead adopted the approach used in PeerReview4All, which optimizes a “leximin fairness” metric (Stelmakh et al., 2021a, code available at https://github.com/niharshah/peerreview4all). This method optimizes the similarity for the most “disadvantaged” proposals; i.e., proposals that have the least overall expertise among reviewers. Once the assignment for the most disadvantaged proposal is optimized, the algorithm moves to the next most disadvantaged, and so on. Proposals in fields with many qualified reviewers are assigned last, as they have more reviewer options available. The metric used by PeerReview4All is solved using the gurobi python optimizer (Gurobi Optimization, LLC, 2023).
@@
The PeerReview4All code was adapted to handle special cases. Student reviewers specify a mentor, who may also serve as a reviewer for other proposals. To ensure fairness, we required that the proposal sets for students and mentors be disjoint. After each round of assignments, the similarity matrix was updated to mark as conflicted any proposals previously assigned to a reviewer or their mentor. In addition, conflicts were updated for reviewers assigned multiple sets, to ensure that they were not assigned the same proposal twice.
@@
To prevent “niche” proposals from being repeatedly reviewed by the same group, further restrictions were imposed on the number of proposals that a reviewer could be assigned within the same topic. These restrictions were specifically applied to proposals related to Very Long Baseline Interferometry, Target of Opportunity, and solar studies.
3 Results
@@
Using the framework described in Section 2 to compute the similarity between proposals and reviewer expertise, we now assess whether the new process improved proposal assignments compared to previous cycles. The new assignment approach was implemented in ALMA Cycle 10, and we evaluate its effectiveness by comparing it to Cycles 8 and 9. To ensure a uniform comparison, we computed similarities for all three cycles using the same LDA model trained on the same set of proposals.
3.1 Evaluating similarity computation on Cycle 8 and Cycle 9 data
@@
We evaluated the effectiveness of the similarity scores (calculated as described in Section 2.2) by using data from previous review cycles. Specifically, we examine whether the computed similarities between a reviewer’s expertise and the proposal topics accurately reflect the reviewer’s expertise. During the distributed peer review process used by ALMA in Cycles 8 and 9, reviewers were asked to self-assess their expertise on each assigned proposal by indicating if the assigned proposal is in their field of expertise (i.e., the review considers themselves an “expert”), they are somewhat knowledgeable, or they have little to no knowledge of the topic. If our similarity scores are a reliable measure of expertise, we would expect that reviewers who self-identify as experts in a proposal topic will, on average, have higher similarity scores compared to those who report having little or no knowledge of the topic.
@@
Figure 3: Cumulative distributions of the similarities between the reviewer’s expertise and the topic of the assigned proposal in Cycle 8 (left) and Cycle 9 (right), when categories and keywords were used to assign proposals. The similarities were computed retroactively to evaluate the machine learning algorithm. The results are shown for proposal assignments where reviewers indicated they are experts (solid curve), had some knowledge of the proposal (dashed), and had little/no knowledge of the proposal (dotted). Similarities tend to be higher when reviewers indicate expertise in the proposal, suggesting that these similarities have predictive value for assessing the suitability of reviewer assignments.
3.2 Comparing similarities for Cycle 9 and Cycle 10 assignments
In this section, we analyze the role of the optimization procedure for computing the assignments (Section 2.3) given the computed similarities. In Cycle 9, proposals were assigned to reviewers based on the proposal category and keywords (Donovan Meyer et al., 2022); i.e., the “old” algorithm. In Cycle 10, we used the similarities and assignment algorithm described in Section 2; i.e., the “new” algorithm. In both cycles, a small percentage of initial assignments was reassigned after the reviewers declared conflicts of interest. [^4] The percentage of assignments where reviewers declared a conflict of interest increased in Cycle 10 to 2.4% from 1.4% in Cycle 9, as more reviewers declared the assignments conflicted scientifically with their own proposal when the machine learning algorithm was adopted.
Figure 4 shows a histogram of the similarity scores for the actual proposal assignments in Cycle 9 (blue histogram) and Cycle 10 (orange histogram), where the latter used the assignment optimization method described in Section 2.3. We find a substantial increase in the similarity between the reviewer’s expertise and the assigned proposals with the new algorithm, as the median similarity increased from 0.20 in Cycle 9 to 0.71 in Cycle 10.
In principle, the increase in similarity scores could be attributed to demographic changes in the reviewer or proposal pools, but this seems unlikely. In Cycle 9, 33% of the reviewers were new compared to Cycle 8, yet the median similarity score remained nearly unchanged (0.18 in Cycle 8 vs. 0.19 in Cycle 9; see Figure 3). Similarly, in Cycle 10, the proportion of new reviewers was again 33% relative to Cycle 9, but the median similarity score increased to 0.71. Given that reviewer turnover remained consistent across cycles and that the increase in similarity coincided with a change in the assignment algorithm, it is more plausible that the improvement reflects an improved alignment between reviewer expertise and proposal topics, rather than demographic changes.
@@ 3.3 Cycle 10 reviewer survey
After completing their reviews, reviewers had the option to complete a survey to self-assess their expertise on their assigned proposals. While self-reported expertise is subjective and may vary with individual perceptions of “expertise” (see, e.g., Karpen, 2018), the survey offers an indication of the suitability of the review assignments. More importantly, trends across cycles reveal how reviewers perceive the quality of their assignments, assuming minimal changes in reviewer demographics between cycles.
Figure 5 shows a histogram of survey results from Cycles 8, 9, and 10. Following the implementation of the new assignment process in Cycle 10, the percentage of reviewers identifying as experts on their assigned proposals increased from 45% in Cycle 9 to 65% in Cycle 10. Simultaneously, the proportion of assignments where reviewers reported little or no knowledge of the proposals dropped from 10% in Cycle 9 to 5% in Cycle 10. These results indicate that the new assignment algorithm achieved a better overall match between reviewers’ expertise and their assigned proposals.
@@ 3.4 Topological accuracy of the proposal rankings
In this section, we evaluate a key tradeoff in expertise-based reviewer assignments: balancing the expertise of assigned reviewers with the topology of the assignments. Strategies to assign proposals to reviewers typically prioritize maximizing relevant scientific expertise. However, strictly focusing on expertise can lead to the isolation of certain proposal groups, especially those within the same subtopic. For example, imagine a scenario where half of the reviewers are exclusively assigned to review one group of proposals, while the other half reviews the rest. This division hampers direct comparison between the two groups, limiting the ability to rank proposals effectively across groups.
@@
To estimate the relative accuracy of the actual assignments in Cycle 10, we set the error in the actual assignments to unity. We then ran simulations where the reviewers were assigned proposals at random across any scientific category, subject to the constraint that each reviewer is assigned 10 proposals and each proposal is assigned to 10 reviewers. The random assignment strategy is known to be optimal (Shah et al., 2016) in terms of topology when all reviewers have identical expertise. We find that the relative uncertainty is 0.86 for the random assignment. Thus, optimizing assignments for expertise increases the uncertainty in final outcomes by only 14% based on the topology of the proposal assignments.
4 Summary and conclusions
The ALMA peer review system has undergone significant evolution in response to the increasing volume of submitted proposals. First, ALMA adopted a distributed peer review system to streamline the logistics of reviewer identification (Donovan Meyer et al., 2022). This paper presents the implementation and outcomes of the next major advancement: the integration of machine learning (Blei et al., 2003) and optimization techniques (Stelmakh et al., 2021a) to assign proposals to reviewers.
The assignment process begins by predicting each reviewer’s expertise for every submitted proposal. This is achieved through topic modeling using machine learning techniques, which analyze the topics of submitted proposals and infer reviewers’ expertise based on their proposal submission history. We calculate the similarity between a reviewer’s expertise and the topics of each proposal, using this metric as a measure of the reviewer’s suitability. To further enhance the accuracy of assignments and address potential mismatches in the machine learning predictions, the process incorporates keywords provided by proposal teams alongside expertise keywords specified by reviewers, ensuring more robust and precise reviewer-proposal matches.
The second step in the assignment process employs optimization algorithms that are used in computer science (Stelmakh et al., 2021a). The optimization algorithms assign proposals to reviewers by prioritizing the most “disadvantaged” proposals (i.e., those with the lowest similarity scores and least available reviewer expertise), thereby improving the overall quality of matches.
Various evaluation metrics highlight the effectiveness of the new algorithm. Testing the similarity computations on proposal assignments from Cycles 8 and 9, which were based solely on keywords, reveals a strong correlation between the similarity scores and reviewers’ self-reported expertise, demonstrating the predictive value of these scores. After applying the new assignment algorithm in Cycle 10, the median similarity between proposal topics and reviewer expertise increased significantly, from 0.20 in Cycle 9 (under the previous algorithm) to 0.71 in Cycle 10. This improvement reflects a substantial enhancement in the alignment of expertise with assignments. The reviewer survey validated these results, showing that the proportion of assignments where reviewers reported high expertise rose from 45% in Cycle 9 to 65% in Cycle 10, while assignments with little or no expertise dropped from 10% to 5%. Moreover, the new system completely eliminated the need for reassignment due to gross mismatches, saving three to five days of manual effort per cycle.
Despite these successes, challenges remain. Since a reviewer’s expertise is inferred from their proposal history, machine learning algorithms can be quite effective in assigning proposals that are similar to a reviewer’s own submitted proposal. This raises questions of potential conflicts of interest (which may or may not be flagged by the reviewer). For instance, in competitive reviewing scenarios, reviewers may downrate competing papers to favor their own (Merrifield & Saari, 2009; Balietti et al., 2016), which may be mitigated using optimization methods to prevent such outcomes (Alon et al., 2011; Dhull et al., 2022) or detecting such behavior (Stelmakh et al., 2021b). Competitive reviewing can also incentivize collusion rings where reviewers manipulate assignment systems to support papers of colluding researchers in exchange for favors (Littman, 2021; Vijaykumar, 2020; Hsieh et al., 2024). Randomization in assignments can mitigate such risks, though not entirely (Jecmen et al., 2020). Addressing these challenges remains an ongoing effort.
While currently employed automated assignment systems maximize each assigned reviewers’ expertise for the papers, it raises questions of whether reviewers should be assigned such a narrow range of topics or if they should be assigned more diverse proposals (Goyal et al., 2024). In addition, the topology of proposal assignments impacts the accuracy of the final rankings, as discussed in Section 3.4. Future work should explore these factors to identify optimal trade-offs.